### v1

# Trust Me, I Can Convince You: The Contextualized Argument Appraisal Framework

Abstract
Emotions, which influence how convincing an argument is, are developed in context of the self and sender, and therefore require modeling the cognitive evaluation process. While binary emotionality has been studied in argument mining, and the cognitive appraisal has been modeled in general emotion analysis, these fields have not been brought together yet. We therefore propose the Contextualized Argument Appraisal Framework that contextualizes the interplay between the sender, receiver, and argument. It includes emotion labels, appraisals, such as argument familiarity, response urgency, and expected effort, as well as convincingness variables. To evaluate the framework and pave the way to computational modeling, we perform a study in a role-playing scenario, mimicking real-world exposure to arguments, asking participants to disclose their emotion, explain the main cause, the argument appraisal, and the perceived convincingness. To consider the subjective nature of such annotations, we also collect demographic data and personality traits of both the participants and the perceived sender of the argument. The analysis of the resulting corpus of 800 arguments, each annotated by 5 participants, reveals that convincingness is positively correlated with positive emotions (e.g., trust) and negatively correlated with negative emotions (e.g., anger). The appraisal variables disclose the importance of the argument familiarity. For most participants, the content of the argument itself is the primary driver of the emotional response.
Trust Me, I Can Convince You:
The Contextualized Argument Appraisal Framework
Lynn Greschner, Sabine Weber,  and Roman Klinger Fundamentals of Natural Language Processing, University of Bamberg, Germany {lynn.greschner,sabine.weber,roman.klinger}@uni-bamberg.de
1 Introduction
Figure 1: Illustration of the Contextualized Argument Appraisal Framework. The perception of the argument, context and sender are cognitively evaluated and assessed as having a negative consequence for the receiver (e.g., losing a job if plastic bottles were banned).
People are frequently exposed to argumentation, where the main goal is persuasion (Habernal and Gurevych, 2016a). Argument quality plays a key role in this process, depending on logical structure (logos), speaker credibility (ethos), emotional appeal (pathos), and contextual relevance (kairos) (Aristotle, 1991; Schiappa and Nordin, 2013). While there is a substantial amount of research on the overall argument quality and convincingness (Habernal and Gurevych, 2016b; Gleize et al., 2019; Wachsmuth et al., 2017; Lauscher et al., 2020), the emotional appeal of arguments remains underexplored in the natural language processing (NLP) community.
One potential reason is that capturing emotions in text is a challenging task due to their contextualized and implicit nature (Casel et al., 2021; Klinger et al., 2018; Koga et al., 2024; Lee and Lau, 2020). This is particularly true in arguments, which rarely directly express an emotion that is intended to be caused. Figure 1 shows an example in our framework. While the argument appears non-emotional, various factors influence the emotional response: The prior beliefs or personality of the receiver, the relationship with and properties of the sender, or the content of the argument itself. Appraisal theories serve as a suitable basis to model the associated cognitive component (Ellsworth and Scherer, 2003). While they received some attention in general emotion analysis (Troiano et al., 2023), there is neither a psychological framework of argument appraisal, nor is there work in NLP on fine-grained modeling of emotions in arguments.
We close this gap by proposing the Contextualized Argument Appraisal Framework which focuses on the subjective perception of arguments. It encompasses emotion categories, a dedicated set of appraisal variables such as argument familiarity, expected effort, or the consequences on the own goals, and the perceived convincingness, all in context of variables of the (perceived) argument sender, the receiver, and the argument text itself. The framework therefore is an emotion argument analysis approach under the paradigm of perspectivism (Frenda et al., 2025).
In this paper, we use the proposed framework and data from the corresponding study to answer the following research questions:
  • RQ1: How do discrete emotions and appraisals correlate with argument convincingness?
  • RQ2: Which contextual factors shape the emotional response to arguments?
Next to answering these questions, the main contributions are the novel framework and the associated corpus ContArgA, consisting of 4000 individual annotations for 800 arguments [^1] The data will be made available to the research community under a Creative Commons license..
2 Related Work
We now summarize relevant research on argument mining, emotion analysis, and appraisal theories to contextualize our work.
2.1 Argument Mining and Convincingness
Previous work has looked in depth at the textual qualities that make an argument convincing (Habernal and Gurevych, 2016a, b). However, the argument itself is not the only driver of the argument’s convincingness; both sender and receiver play a part. Al Khatib et al. (2020) and Lukin et al. (2017) examine the personality of the receiver using the Big Five personality traits, finding that certain personality types are more receptive to certain types of arguments. Lukin et al. (2017) also show that the prior belief is important, with receivers who hold only weak beliefs about the argument topic being more easily convinced. Durmus and Cardie (2018) show that the prior belief of the receiver plays a more important role for the prediction of convincingness than linguistic features of the argument. In line with that, Rescala et al. (2024) demonstrate that LLMs perform on par with humans when predicting which arguments would be convincing for individuals with specific demographics and beliefs.
We follow this work by collecting not only assessments of argument properties, but also properties of the annotators (age, gender, education level and Big Five personality traits). We specifically collect the stance of the receiver toward the argument topic beforehand and check for stance change after an argument was displayed. Additionally, we collect information about the sender that annotators imagine after reading the argument (age, education level, personality traits and a free text description).
2.2 Emotion Analysis and Appraisals
There are two main types of emotion models for emotion analysis: categorical and dimensional. Ekman’s basic emotion model, as an example for categorical models, consists of six discrete emotions: anger, surprise, disgust, joy, fear, and sadness (Ekman, 1992). Dimensional models represent emotions along continuous axes in a multidimensional space. The most prominent dimensional model in NLP is the Circumplex Model of Affect (Posner et al., 2005), evaluating emotions in terms of valence and arousal. Appraisal theories also constitute a dimensional model, in which emotions are considered in terms of the cognitive evaluation (appraisal) of an event (Scherer et al., 2001). While multiple frameworks of appraisal theories exist (Roseman, 1984; Roseman and Smith, 2001; Scherer, 2009), there is no consensus on the concrete set, leading to specialized framework for comparably niche areas, such as conspiracy theories (Pummerer et al., 2024). Common variables in appraisal frameworks do, however, encompass aspects of agency, pleasantness, consequences on the self, responsibility, or expected effort and novelty.
While most work on emotion analysis is on categorical models, there is considerable work that employs appraisal theories (Klinger, 2023), including work on coping strategies (Troiano et al., 2024), social media analysis (Stranisci et al., 2022), or emotion event self reports (Hofmann et al., 2020). The largest set of appraisal variables for emotion analysis has been proposed by Troiano et al. (2023), which we use as inspiration in our work. There is no appraisal framework for arguments in psychological research (but the recognition that it would be important, Dillard and Peck (2000)). Therefore, our contribution is both on the conceptual side and the empirical natural language processing side.
Dimension Description Values
Suddenness the argument appears sudden or abrupt to the receiver 1–5
Suppression the receiver tries to shut the argument out of their mind 1–5
Familiarity the argument is familiar to the receiver 1–5
Pleasantness the argument is pleasant for the receiver 1–5
Unpleasantness the argument is unpleasant for the receiver 1–5
Consequencial Importance the argument has important consequences for the receiver 1–5
Positive Consequentiality the argument has positive consequences for the receiver 1–5
Negative Consequentiality the argument has negative consequences for the receiver 1–5
Consequence Manageability the receiver can easily live with the unavoidable consequences of the argument 1–5
Internal Check the consequences of the argument clash with the receiver’s standards and ideals 1–5
External Check the consequences of the argument violate laws or socially accepted norms 1–5
Response urgency the receiver urges to immediately respond to the argument 1–5
Cognitive Effort processing the argument requires a great deal of energy of the receiver 1–5
Argument Internal Check statements in the argument clash with the receiver’s standards and ideals 1–5
Argument External Check statements in the argument violate laws or socially accepted norms 1–5
Table 1: Appraisal dimensions and descriptions used in our study for argument appraisal.
2.3 Emotions in Arguments
A substantial number of psychological studies point out the role of cognitive argument evaluations for convincingness (Bohner et al., 1992; Petty et al., 1993; Pfau et al., 2006; Worth and Mackie, 1987; Benlamine et al., 2015). Particularly, pathos plays an important role in stance changes (Benlamine et al., 2017). In natural language processing, research on emotions in arguments is limited, where emotional appeal is one factor of many (Habernal and Gurevych, 2016a) or is considered fallacious (Evgrafova et al., 2024).
The convincingness and emotion are related, though studies with large language models show the challenging nature of automatically detecting these properties (Chen and Eger, 2025). There is very little work that studies not only binary emotionality but differentiates emotion categories. Greschner and Klinger (2025), as one of the few cases, demonstrate that the discrete emotion category enhances LLM predictions for emotionality in arguments. Habernal and Gurevych (2016a, b) constructed and analyzed a corpus for convincingness strategies, including an emotion layer. In line with that, Wachsmuth et al. (2017) find a weak positive correlation between the emotional appeal of arguments and convincingness when analyzing 15 dimensions of arguments. Lukin et al. (2017) focus on audience factors and argument convincingness and find that information about personality traits of the audience improves the automatic prediction of belief change for emotional arguments.
Our work considers sender and receiver properties in the context of argument convincingness, focusing on the implicitly evoked emotions in receivers. To the best of our knowledge, this work is the first to introduce appraisal theories to arguments in NLP.
3 Contextualized Argument Appraisal Framework
We illustrate the Contextualized Argument Appraisal Framework (CAAF) in Figure 1. CAAF considers arguments within a situational context with the sender, world knowledge, and the self (receiver). The receiver cognitively evaluates, in a synchronized but not necessarily sequential manner, the appraisal, emotion category, and convincingness of the argument.
3.1 Emotions
The cognitive appraisal of the argument leads to a subjective experience that corresponds to a discrete emotion category. This emotion category, along with the perceived emotion intensity, are subjective variables of the receiver and need to be assessed by them (i.e., cannot be externally reconstructed by annotators).
Arguments are situated within context and presented by a sender. We include these factors in CAAF to study the major source of the emotion stimulus: how is the emotion evoked and why is the emotion evoked.
3.2 Appraisals
To reconstruct the subjective cognitive evaluation process of a presented argument in context, we introduce appraisal variables for arguments, the main and novel component in our framework. We base the appraisals used in our study on the work by Troiano et al. (2023), adapting a subset (suddenness, familiarity, pleasantness, unpleasantness, internal check, external check) to match the appraisal of arguments, as described in Table 1. Furthermore, we design nine additional appraisal dimensions: suppression, consequencial importance, positive consequentiality, negative consequentiality, consequence manageability, response urgency, cognitive effort, argument internal check, and argument external check.
Appraisal dimensions allow us to capture the underlying cognitive evaluations that give rise to emotional responses. They enable a systematic study of subjective argument properties.
3.3 Convincingness
Central to the CAAF is the convincingness assessment of arguments. Convincingness refers to the quality of an argument that leads its audience to believe the content of the argument is factually true or morally right. In our framework, we acquire this variable by a self-assessment of the argument receiver, but, in addition, obtain the psychologically more valid property of an observed stance change. We do, however, use the convincingness variable in our studies due to pragmatic design choices in the data collection described later.
3.4 Context
The variables above can only be assessed in the situational context, which therefore requires information about the sender and the receiver. Text is rarely produced in a contextual vacuum. Rather it is influenced by demographic properties, personality traits and preconceived notions of the people involved. Capturing these features allows us to create models that take individual nuances into account, both in the ground-truth properties of the receiver (e.g., their education level) and the perceived properties of the sender (e.g., does the argument sound like it was said by an educated person).
Figure 2: Flowchart of the annotation workflow. Yellow fields indicate information collected about the receiver (the annotator), green fields indicate argument and scenario display and blue fields indicate argument specific annotation.
4 Corpus Creation
We now describe our annotation setup based on CAAF, to study the correlations of emotion, appraisal, and convincingness, and pave the way for computational modeling. Screenshots of the study are provided in Appendix E.
4.1 Overview
Figure 2 displays the flowchart of our annotation setup. The study begins with assessing the participant’s emotional state, personality, and stance toward the topic. Subsequently, participants are immersed in a role-playing scenario simulating a town-hall meeting, designed to approximate real-world argumentative contexts and to evoke authentic emotional and cognitive engagement. Exact scenario descriptions and the involvement questions are Appendix D. The role-playing approach, used in NLP to create anntotated corpora Troiano et al. (2024), puts individuals in the shoes of fictional characters.
We ensure emotional involvement in the role-playing scenario, participants describe where they would sit and what they would see in their surroundings. Next, a speaker presents an argument, which participants annotate. They then generate an argument themselves to increase involvement (this data is not evaluated in this paper). Participants further answer demographic and personality trait questions about the speaker. This process is repeated for a second argument within the same role-playing scenario. Lastly, the participants disclose own demographics. Seven attention checks are inlcuded to ensure high data quality, see Appendix E for details.
4.2 Variables
We now turn to a detailed description of the annotated variables.
4.2.1 Argument
The following section describes the collected variables on the argument level in detail. All argument variables are annotated with the subjective perception of the argument by the receiver.
Emotions.
Our emotion label set comprises Ekman’s basic emotions anger, disgust, fear, joy, sadness, surprise, expanded by shame, guilt, pride, (Greschner and Klinger, 2025), trust and relief (Troiano et al., 2023). Each emotion is rated with its intensity on a 5-point Likert scale. We refer to the emotion with the highest rating as the dominant emotion. In the case of ties participants are requested to chose one. We assess the perceived interest and boredom of the argument on a 5-point Likert scale at a different point in the study and with a different formulation than the emotion questions, see Appendix E for the formulations.
Emotion Reason.
Participants are instructed to provide a short textual explanation of the evoked emotions using a free text field without length restriction.
Emotion Influence.
Participants answer three questions about the influence of the sender, argument content, and receiver on a 5-point Likert scale. The questions are phrased in the format “[The speaker identity, the argument itself, who I am] was important for the emotion I developed”.
Appraisals.
We phrase each appraisal dimension as an affirmation, for instance, for response urgency we formulate: “I feel the urge to immediately respond to the argument”. Every appraisal dimension is evaluated on a 5-point Likert scale (not at all, …, extremely).
Manipulation.
For each argument, participants provide their perceived degree of manipulative intent. We formulate the question: “Do you feel manipulated by this argument?” and allow answers on a 5-point Likert scale (not at all, …, extremely) [^2] In preliminary studies, we found perceived manipulation to decrease convincingness. In ContArgA, there is a weak, significant negative correlation between manipulation and convincingness, which we leave for future work for a more in-depth analysis..
Convincingness.
Similar to the manipulation, participants annotate the perceived convincingness for the question “How convincing is this argument for you?” on a 5-point Likert scale.
4.2.2 Receiver
Given that individual differences among receivers can affect the perceived persuasiveness of arguments, as demonstrated by Lukin et al. (2017), our study accounts for these variations by employing the following measurements.
Stance in Relation to Argument Topic.
Participants are asked to provide their stance towards a the topic of the displayed town hall discussion on a 5-point scale (strongly disagree, …, strongly agree). This question is asked once prior to argument exposure and once after each argument, allowing us to measure potential change in the stance after specific arguments. Additionally, participants provide their familiarity with the topic on a 5-point scale (not to be confounded with the appraisal of familiarity with the argument).
Demographics.
Previous work on argument quality has shown that the lack of information about the annotators of corpora poses an impediment to perspectivist approaches to the topic (Romberg et al., 2025). To avoid this shortcoming, we collect the age, gender and education level of the participants.
Personality Traits.
Lukin et al. (2017) demonstrate that more conscientious people are rather convinced by emotional arguments. Since we hypothesize that evoked emotions are one of the primary drivers of the argument’s convincingness, we use the Ten-Item Personality Inventory questionnaire (Gosling et al., 2003) to assess the extraversion, agreeableness, conscientiousness, emotional stability, and openness to experiences of the participants.
Argument Topic Emotion–Convincingness Labels
“people simply can’t compete with others in the same professional setting once they reach a certain age, mandatory retirement is necessary for fair competition in a limited job market” Retirement guilt–4; disgust–1; relief–3; sadness–2; surprise–1
“women in polygamous marriages have been shown to be at much greater risk of abuse than those in monogamous marriages” Polygamy disgust–3; sadness–2; anger–3; sadness–3; surprise–2
“there are direct links to drinking bottled water and certain types of disease, even cancer” Plastic bottles fear–4; trust–1; anger–1; guilt–2; surprise–2
Table 2: Examples from the ContArgA corpus from various topics. The emotion & convincingness column contains five individual emotion and convincingness labels for each argument. The labels from different annotators are indicated as emotion–convincingness score.
4.2.3 Sender
In our study setup, the age (old/young) and gender (male/female) of the argument-presenting sender are systematically manipulated. While we are aware that neither age nor gender are binary variables, this simplified setup allows us to investigate whether these properties influence persuasiveness more easily. Each argument is attributed to one single sender, whose age and gender are assigned according to a balanced randomization scheme across topics to ensure approximately equal representation of all four demographic combinations. We present participants with a description of the sender and the argument in the format “A [speaker age] [speaker gender] approaches the microphone and makes a statement: [argument]”.
Following the argument exposure, participants answer the same demographic and TIPI questionnaire questions described in 4.2.2 about their perception of the sender. This approach allows us to examine the interplay between speaker characteristics and audience perception within argumentative discourse. Additionally, the participants can give details about the imagined sender in a free text field. We leave a more in-depth analysis of these variables for future work.
4.3 Crowd-sourcing Details
We use the platform Prolific [^3] https://www.prolific.com/ with a self-implemented study framework based on Streamlit [^4] https://docs.streamlit.io/. Participants are required to live in the United Kingdom or Ireland, have English as their first, native, and primary language, and have an approval rate of 95–100%. Each participant answers the survey for two arguments. We pay each participant 3.48 £ for one survey, which on average takes 22.5 minutes. The study includes 7 attention checks. Participants can participate up to 40 times, and therefore annotate up to 80 arguments. In total, the cost of the study amounts to 9,404 £ . The contributing participants in our studies were on average 40.4 years old (18 minimum, 81 maximum). From this set, 1,715 identified as female, 1,541 as male, 374 as non-binary, and 30 as genderfluid/non-conforming. Further, 14 specified their own gender, 12 preferred not to answer the question, and 8 as questioning [^5] Note that on the Prolific platform, participants cannot provide their gender in such detail. We distributed our study as 40% female, 40% male, 20% non-binary and obtained the gender variable in the study..
4.4 Datasets
The UKPConvArgv1 (Habernal and Gurevych, 2016b) and IBM-Rank-30k (Gretz et al., 2019) corpora, hereafter UKP and IBM, serve as sources for the arguments. Both datasets provide isolated arguments and their respective discussion topic. Isolated arguments allow us to investigate the initial argument appraisal without adding noise due to lengthy debates containing multiple arguments. We sample 14 and 26 topics from UKP and IBM, respectively (see Appendix A). For each topic, an equal number of pro and con arguments is sampled. Arguments are manually annotated for comprehensiveness, persuasive attempt, and stance alignment. Importantly, we exclude arguments that point toward the age or gender of the sender, since we manipulate these variables in our study (see Section 4.2.3). See Appendix A for more details on the manual filtering. We exclude 170 annotations due to failed attention checks and randomly select 800 arguments from the 953 valid ones for our study.
5 Data Analysis
Our study results in a corpus of 4,000 individual argument annotations of 800 arguments from 39 topics with 5 individual annotations. Details about the topics can be found in Appendix C.
Table 2 displays example arguments with emotion and convincingness labels. The average length of the arguments is 32.1 tokens, 1.7 sentences. The average value of convincingness is 2.7 (min=1, max=5); the values also depend on the topic, with retirement being most (3.2) and atheism being least convincing (2.0). The frequencies of the dominant emotions are, in descending order: trust (709), surprise (564), sadness (561), anger (521), relief (382), disgust (292), pride (198), joy (193), shame (100), fear (90), guilt (84). The average emotion intensity value across emotions is 1.61, with trust showing the highest (2.02), guilt showing the lowest (1.27) value.
We now turn to the example arguments from ContArgA, displayed in Table 2, in more detail. The first argument evokes 5 different emotions and convincingness labels from 1–4. In contrast, the second argument evokes negative emotions (disgust, sadness, anger) with similar convincingness values, with the exception of surprise.
The second argument discusses the greater risk of abuse of women in polygamous marriages, intuitively evoking negative emotions. Surprise may arise if the receiver is unfamiliar with the topic. For the third argument, we again find negative emotions (fear, anger, guilt). Fear may result from the risk of developing cancer, while a person regularly using plastic bottles could feel guilty. The argument evoking trust is less intuitively clear. These examples illustrate the wide variety of emotions evoked in different participants by the same arguments.
Inter-Annotator Agreement.
The contextualized argument appraisal framework is designed to capture individual argument appraisals on divisive topics. We intentionally construct it to elicit differences between receivers of arguments rather than to derive one ‘gold-label’. As expected, the inter-annotator agreement on single variables is therefore low, with a Krippendorff’s Alpha of 0.06 for convincingness and Fleiss’ Kappa of 0.17 for the most dominant emotion category of an argument. These numbers do not suggest a low quality of the data, but justify the need for modeling and analysis under the paradigm of perspectivism.
Figure 3: Emotion Intensities and Convincingness.
Variable r p
Anger -0.222 0.000
Disgust -0.233 0.000
Fear 0.006 1.000
Guilt 0.113 0.000
Joy 0.447 0.000
Pride 0.465 0.000
Relief 0.510 0.000
Sadness -0.125 0.000
Shame -0.052 0.027
Surprise -0.085 0.000
Trust 0.578 0.000
Manipulation -0.104 0.000
Interest 0.691 0.000
Boredom -0.491 0.000
Suddenness -0.344 0.000
Suppression -0.296 0.000
Familiarity 0.336 0.000
Pleasantness 0.576 0.000
Unpleasantness -0.363 0.000
Consequential Importance 0.151 0.000
Positive Consequentiality 0.400 0.000
Negative Consequentiality 0.217 0.000
Consequence Manageability -0.018 1.000
Internal Check -0.105 0.000
External Check -0.366 0.000
Response Urgency -0.249 0.000
Cognitive Effort -0.078 0.000
Argument Internal Check -0.124 0.000
Argument External Check -0.498 0.000
Table 3: Pearson correlations ($r$) and corresponding $p$ values (with Bonferroni correction) for all variables and convincingness.
6 Results
6.1 RQ1: How do discrete emotions and appraisals correlate with argument convincingness?
We analyze ContArgA to investigate the interplay between evoked emotions and convincingness.
Emotions.
We aim at understanding the specific emotions that are evoked in the receivers of arguments to examine their interplay with perceived convincingness. To this end, we analyze ContArgA arguments for relationships between discrete emotion categories and convincingness labels.
Table 3 displays the Pearson’s correlation values between all variables and convincingness. All correlations between emotions and convincingness are significant, except for fear. The strongest positive effects are with trust (.58), relief (.51), pride (.47), and joy (.45). In contrast, anger (-.22) and disgust (-.23) show negative correlations with convincingness. These findings align with Greschner and Klinger (2025).
Emotion Intensities.
Figure 3 allows a closer look at these correlations, based on intensity values. For trust, the emotion intensity and convincingness have an almost linear relationship, similarly for relief, pride, joy. For fear, shame, and guilt, the emotion intensity stays relatively stable across convincingness values. In contrast, for anger and disgust, the highest intensity values are found for the lowest convincingness values, with a drop at value 2 (little convincing), continuing in a moderate intensity (around 1.4) across higher convincingness values. Similar, yet weaker trends are found for sadness and surprise, with on average higher emotion intensity values (around 1.7).
Figure 4: Analysis of discrete emotion categories and appraisal dimensions in arguments. Each row is an appraisal, each column is an emotion, and each cell is the average appraisal value for that emotion.
Appraisals.
We now focus on the analysis of emotions and appraisals in arguments, illustrated in Figure 4. For each emotion, the average appraisal value is displayed in the heatmap. The values are in line with other applications of appraisal theories to emotion analysis in natural language processing (Troiano et al., 2023, Figure 8), confirming the validity of the approach.
Unpleasantness, internal check and external check appraisals consistently show the highest ratings for anger (3.0, 2.8, 2.7) and disgust (3.2, 2.8, 2.9). Intuitively, arguments evoking negative emotions are being appraised as unpleasant, clashing with the receiver’s ideals, and violating social norms. In contrast, joy, pride, and relief show high values for pleasantness (3.5, 3.2, 3.0), familiarity (3.1, 3.3, 3.2), and positive consequentiality (2.5, 2.5, 2.5). This indicates that arguments being appraised as beneficial, familiar, and rewarding evoke positive emotions.
Beyond categorical emotions, appraisals also explain convincingness: we find significant correlations for all appraisal dimensions (see Table 3) except consequence manageability. Pleasantness shows a strong positive (.58), positive consequentiality a moderate (.40), and familiarity a moderate (.34) positive correlation with convincingness. In contrast, we find strong negative correlations with argument external check (-.50), unpleasantness (-.36), and external check (-.36).
Both discrete emotion and appraisal analyses demonstrate significant correlations between positive emotions and convincingness, and negative correlations between negative emotions and convincingness. These results align with Greschner and Klinger (2025), but contradict that fear increases convincingness (Dillard and Anderson, 2004).
Emotion Sender Argument Receiver
Anger 1.86 3.78 3.23
Disgust 1.82 3.72 3.25
Fear 1.78 3.68 3.11
Guilt 1.65 3.27 2.81
Joy 2.13 3.73 3.21
Pride 2.24 3.79 3.24
Relief 1.98 3.70 3.14
Sadness 1.84 3.41 3.09
Shame 2.00 3.34 2.86
Surprise 1.86 3.39 2.69
Trust 1.82 3.47 2.97
Avg. 1.89 3.56 3.05
Table 4: Average values of how important the sender, the argument, and the receiver were for the emotional response of the receiver. The importance of the the 3 dimensions was rated on a 1–5 scale.
6.2 RQ2: Which contextual factors shape the emotional response to arguments?
After analyzing which emotions are developed and how they are correlate with the argument’s convincingness, we now turn to understanding how these emotions are developed. We analyze the participant’s (argument receiver’s) textual explanations about the evoked emotions and the strongest influence on the emotion (sender, argument, receiver). We also analyze how strong (1–5) the influence (sender, argument, receiver) was for the development of the evoked emotion (see Table 4). The argument (3.6) is the primary driver of the emotional response, followed by the receiver themselves (3.1) and the sender (1.9). Receivers rarely attribute guilt and shame to being evoked by the sender.
Pride is often attributed to the argument itself, which is also the most frequent emotion and has the highest correlation to convincingness. Similar patterns are found for anger, which has however has the strongest negative convincingness correlation. In summary, the argument itself is the primary driver of the emotional response.
7 Conclusion
In this paper, we introduce the Contextualized Argument Appraisal Framework (CAAF), an approach to explain how arguments are cognitively appraised. In a synchronized process, individuals also develop an emotion and a perceive a particular convincingness. The analysis of the resulting corpus ContArgA, the data resulting from a role-playing annotation study based on CAAF, shows that appraisals are appropriate for explaining argument convincingness. We release the data, enabling research on arguments with focus on contextual (sender, receiver, argument) properties.
Our work demonstrates that affective responses are not merely superficial reactions but are integrally tied to how arguments are processed. Future work shall study CAAF from a computational perspective. We hypothesize that the integration of appraisal and categorical emotion variables improves the performance of automatic convincingness estimates and make the decision process transparent. Further, models that automatically assess the subjective appraisal and emotion are valuable in itself for corpus and argumentation studies.
Particularly, CAAF helps to disentangle the influences that shape a person’s reaction to arguments, because it captures the role of properties of the sender, receiver, and the argument itself. In downstream systems, these factors can contribute to argument mining and content moderation systems, for example to support studies on desinformation and argumentation strategies. The framework lays the groundwork for context-sensitive and emotionally aware computational systems that reflect how people engage with persuasive text.
Limitations
One of the limitations of this work is the focus on the English language and a relatively limited geographical spread of the annotators involved in the creation of our corpus. We make this choice to ensure higher data quality, limiting a possible mismatch between dialects or regional language usage. On the other hand, a wider linguistic and geographic spread would allow us to make statements about the nature of emotions in arguments that are more comprehensive. While we see these shortcomings in the annotation itself, the Contextualized Argument Appraisal Framework stands independent of its implementation in this specific annotation scenario and can be used across various languages.
Previous work has also examined multi-turn argumentative conversations. While this allows an observation of how emotions and appraisals develop over the course of a longer exchange, this is out of the scope of our current setting of the framework. Adapting it for multi-turn interactions is another avenue for future work.
Regarding our study, each participant annotates two arguments from one topic. The argument selection from the topic is randomized, which can lead to biases due to the prior stance of the participant and the order of pro and con arguments. We acknowledge that future work can take a deep dive into the role of argument presentation according to stances, since we report the exact argument in displayed order within our available ContArgA corpus. Further, we allow participation up to 39 times, which ensures that one participant can only annotate two arguments per topic, however, there might be annotators that do in fact participate in all 39 studies (i.e, contributing 78 individual argument annotations out of the 4,000 arguments in ContArgA), therefore introducing annotator bias.
Ethical Considerations
The study is part of a larger project that has been approved, including a description of the presented study, by the ethics board of the University of [Anonymous]. Following recommendations with respect to ethical challenges in emotion analysis (Mohammad, 2022), we point out the following facts about our work: The purpose of the CAAF is to study the interplay of emotions in argument texts while taking contextual factors into account, focusing on the emotions evoked in receivers. We choose categorical emotion labels from emotion theories in accordance with previous work, and expand appraisal labels from dimensional appraisal theories to be applicable to arguments.
We employed crowd-sourcing to collect human annotations for the cognitive appraisal process of arguments. For a given argument, participants provide emotional responses, convincingness, stances toward and familiarity with the topic. In addition, participants provide demographic information, personality traits, and their emotional state before starting the study. We informed the participants that the intent of the study is research with the goal of a scientific publication. We obtained consent from all participants to use their data to analyze emotions and convincingness in arguments as well as publishing their annotation in anonymized from. While there is one free text field in our study, we manually check all of these responses for any hints that would allow tracing back the identity of the annotator, for instance, if a participant unconsciously disclosed personal information. We explicitly make annotators aware that they will not be compensated if they fail more than 2 attention checks. Details about the payment are described in Section 4.3. We are aware that being exposed to arguments can be upsetting. We manually select a comprehensive set of trigger words and warn the study participants about the potentially upsetting nature of arguments. The participants can drop out of the study at any point without consequences.
Acknowledgments
This project has been conducted as part of the Emcona (The Interplay of Emotions and Convincingness in Arguments) project, which is funded by the German Research Foundation (DFG, project KL2869/12–1, project number 516512112). We thank Meike Bauer for her thorough work and the fruitful discussions.
Appendix A Manual Data Filtering of UKP and IBM.
The UKPConvArgv1 (Habernal and Gurevych, 2016b) and IBM-Rank-30k (Gretz et al., 2019) corpora, further referred to as UKP and IBM, serve as the source for the arguments. We aimed for 1000 arguments in total. We approximate needing 40 various topics to capture topical differences; hence, we select all 16 topics from UKP and randomly select 24 topics from IBM. For each topic, we randomly sample 14 arguments per stance (pro/con). For this set, we perform standard data cleaning and pre-processing steps.
We manually check if the argument matches our definition of arguments: The argument must provide reasons supporting a specific claim (standpoint or point of view). The argument must be complete and comprehensive in itself. The reasons must at least be attempted to be logical and reasonable. The argument must follow the goal of persuasion (convince others of the supported standpoint) and must be logical in itself. Further, we exclude arguments using the following criteria. (1) The argument is not in line with our definition of an argument. (2) The argument shows hints toward author characteristics (e.g., “As a woman, I would argue that…”). (3) The argument is a duplicate. (4) The argument does not have a clear stance toward the topic.
For the resulting arguments, we check the stance of each argument for correctness, re-annotate the stance if necessary, and normalize all stances to be either “yes” if the argument supports the topic or “no” if the argument opposes the topic. We normalize the arguments’ topics according to this structure: “We should [verb] [object].” For instance, we normalize the topic “gay-marriage-right-or-wrong” to “We should allow gay marriage”.
This manual effort results in 953 arguments, from which we randomly select 800 for the annotation.
Appendix B Emotion Influence
In Section 6.2 we analyze the importance of the sender, argument, and receiver in detail to understand the drivers of emotional responses. We analyze how important (1-5) the influence (sender, argument, receiver) was for the development of the evoked emotion, individually for each topic and display the results in Table 5. While the main results are discussed and contextualized in Section 6.2, we display the results for all topics here.
Topic Sender Argument Receiver
Weapons 1.67 4.08 3.30
Farming 1.48 3.76 3.37
Zoos 1.51 3.71 3.05
Libertarianism 1.54 3.69 3.18
Gay Marriage 1.85 3.97 3.71
Guantanamo 1.65 3.76 3.16
Cosmetic Surgery 1.78 3.73 2.89
Exec. Compensation 1.64 3.51 3.02
Gender 2.04 3.87 3.30
Naturopathy 1.73 3.56 2.93
Plastic Bottles 1.82 3.63 2.67
TV Books 1.66 3.47 2.62
Int. Property 1.69 3.48 2.66
Entrapment 1.79 3.56 3.34
Personal Pursuit 1.80 3.55 2.93
ZTP Schools 1.76 3.46 3.04
PE Schools 1.74 3.43 3.12
Atheism 1.78 3.47 3.00
Journalism 1.76 3.43 2.81
Holocaust 1.82 3.48 2.96
Economic Sanctions 1.82 3.46 3.06
Prostitution 2.09 3.72 3.24
Spouse 2.03 3.66 3.45
Creationism 1.56 3.17 3.10
Education 1.98 3.57 3.07
Spanking 2.04 3.61 3.33
Polygamy 1.90 3.44 2.98
India 1.92 3.44 2.98
School Uniform 2.00 3.51 2.78
Abortion 2.35 3.85 3.00
Women Combat 2.31 3.77 3.39
Student Loans 1.93 3.36 2.82
Judicial Activism 2.08 3.50 2.83
Fastfood 1.95 3.36 3.09
Father 2.23 3.44 3.20
Retirement 2.38 3.56 2.84
Subsidize Dads 2.18 3.35 2.98
Porn 1.92 2.83 3.08
Firefox 2.22 3.14 2.56
Table 5: Average emotion influence scores across topics.
Appendix C Topics
ContArgA contains 39 distinct topics. We display the short names for the topics, along with a description description of each topic, the number of pro, con, and total arguments in Table 6.
Name Description Corpus Pro Con Total
Abortion We should legalize abortion IBM 9 9 18
Atheism We should adopt atheism UKP 7 7 14
Cosmetic Surgery We should ban cosmetic surgery for minors IBM 7 7 14
Creationism We should believe in creationism UKP 13 13 26
Economic Sanctions We should end the use of economic sanctions IBM 11 11 22
Education We should subsidize vocational education IBM 13 13 26
Entrapment Entrapment should be legalized IBM 13 13 26
Exec. Compensation We should limit executive compensation IBM 13 13 26
Farming We should ban factory farming IBM 11 11 22
Fastfood We should ban fast food IBM 14 14 28
Father It is better to have a lousy father than to be fatherless UKP 13 13 26
Firefox We should choose firefox over internet explorer UKP 11 11 22
Gay Marriage We should allow gay marriage UKP 4 4 8
Gender We should adopt gender-neutral language IBM 12 12 24
Guantanamo We should close Guantanamo Bay detention camp IBM 14 14 28
Holocaust Holocaust denial should be a criminal offence IBM 13 13 26
India India has the potential to lead the world UKP 6 6 12
Int. Property We should abolish intellectual property rights IBM 12 12 24
Journalism We should subsidize journalism IBM 12 12 24
Judicial Activism We should limit judicial activism IBM 13 13 26
Libertarianism We should adopt libertarianism IBM 12 12 14
Naturopathy We should ban naturopathy IBM 14 14 28
PE Schools Physical education should be mandatory in schools UKP 9 9 18
Personal Pursuit We should prioritize the personal pursuit over advancing the common good UKP 13 13 26
Plastic Bottles We should ban plastic water bottles UKP 6 6 12
Polygamy We should legalize polygamy IBM 13 13 26
Porn Porn is wrong UKP 8 8 16
Prostitution We should legalize prostitution IBM 14 14 28
Retirement We should end mandatory retirement IBM 13 13 26
School Uniform The school uniform is a good idea UKP 9 9 18
Spanking We should allow spanking of children UKP 10 10 20
Spouse One should turn in their spouse if they committed murder UKP 12 12 24
Student Loans We should subsidize student loans IBM 13 13 26
Subsidize Dads We should subsidize stay-at-home dads IBM 13 13 26
TV Books TV is better than books UKP 7 7 14
Weapons We should fight for the abolition of nuclear weapons IBM 11 11 22
Women Combat We should prohibit women in combat IBM 12 12 24
ZTP Schools We should adopt a zero-tolerance policy in schools IBM 13 13 26
Zoos We should abolish zoos IBM 13 13 26
Table 6: List of topics in ContArgA, along with original data source (IBM or UKP), pro, con, and total number of arguments.
Appendix D Role-Playing Scenario
To mimic the exposure to an argument in the wild and to increase emotional engagement, we create the role-playing scenario of a town hall meeting that the annotator is part of. The exact role-playing scenario is described in Figure 5. The involvement questions that help the annotator engage with the argument can be seen in Figure 6.
Figure 5: Description of the role-playing scenario in our annotation study.
Figure 6: Involvement question for the role-playing scenario in our annotation study.
Appendix E Study Design
The following section shows the screenshots of out study. After Figure 26 the study loops back to Figure 14 to display the second argument.
Figure 7: Screenshot of our Study, Page No. 1
Figure 8: Screenshot of our Study, Page No. 2
Figure 9: Screenshot of our Study, Page No. 3
Figure 10: Screenshot of our Study, Page No. 4
Figure 11: Screenshot of our Study, Page No. 5
Figure 12: Screenshot of our Study, Page No. 6
Figure 13: Screenshot of our Study, Page No. 7
Figure 14: Screenshot of our Study, Page No. 8
Figure 15: Screenshot of our Study, Page No. 9
Figure 16: Screenshot of our Study, Page No. 10
Figure 17: Screenshot of our Study, Page No. 11
Figure 18: Screenshot of our Study, Page No. 12
Figure 19: Screenshot of our Study, Page No. 13
Figure 20: Screenshot of our Study, Page No. 14
Figure 21: Screenshot of our Study, Page No. 15
Figure 22: Screenshot of our Study, Page No. 16
Figure 23: Screenshot of our Study, Page No. 17
Figure 24: Screenshot of our Study, Page No. 18
Figure 25: Screenshot of our Study, Page No. 19
Figure 26: Screenshot of our Study, Page No. 20
Figure 27: Screenshot of our Study, Page No. 21
Figure 28: Screenshot of our Study, Page No. 22
Figure 29: Screenshot of our Study, Page No. 23
Figure 30: Screenshot of our Study, Page No. 24
Figure 31: Screenshot of our Study, Page No. 25
Figure 32: Screenshot of our Study, Page No. 26
Figure 33: Screenshot of our Study, Page No. 27
Figure 34: Screenshot of our Study, Page No. 28
Figure 35: Screenshot of our Study, Page No. 29